OpenStudy (anonymous):

Determine whether each formula is explicit or recursive. Then find the first five terms of each sequence. a n= 2a n-1 +3, where a1 = 3

OpenStudy (amistre64):

anything that is defined in terms of prior terms is recurrsive

OpenStudy (amistre64):

a1 = 3 a2 = 2(a1) + 3 a3 = 2(a2) + 3 a4 = 2(a3) + 3

OpenStudy (anonymous):

thanks